shift report

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Sat Sep 15 2001 - 19:40:25 EDT

  • Next message: Apache: "Shift report 20010915 08:00-16:00"

    This shift report was not send since it was done from pii4 -
    sorry the belated mailing.
    
    
    Date: Sat Sep 15 16:24:26 2001 Shift: 08:00-16:00
    Personel: fv and fv
    Log book: 8 Pages: 133 - 136
    
    
    I lost whta was written once - so I included the complete
    detailed log that was kept during the run in a editor.
    
    Beam 8-10:15 ,, 14:15-16:00  4 hours
    
    8:41 - during run 4740
         T5Map  problem and correspondinf sync error.
    
    8:42 Ran the Scaler.C program to check current deadtime
         it is 86% live. CLose to vaue reported by DAQ.
    
    8:45  The Brahsm ZDC rate are only 1/3of Phobos/Star while half of Phenix.
          Will remedy next store.
    
    8:50  Blue   Beam ZDC Singles: 3.1K  Ions == 16.5 10**9
          Yellow Beam ZDC Singles: 5.1K  Ions == 15.5 10**9
    
    9:00  The Scaler reading program that calculates the rates (reading the
          Brahms scalers seem to record large rates than the RHIC scaler. This
          is peculiar since it is supposed to be the same signals.
          This is becaues I looked at one - not 4..
    
    9:05  Started/stopped run to adjust scaler rates.
          
    
    9:05  START run 4741.
    
    9:17  checking of Monitor plots from previous run 4740.
    
          Comparing Plots of T3 following the modification of Bob.
    
          Improvements: 
          T3-1-X3 looks similar to T3-1-Y1 
          T3-2-V2 looks smooth across wirenumbers now.
          T3-3-X1,X2,X3 got rid of excessive rates in wire 8-24.
          T3-3-U2 looks good
          T3-3-V2 seem better for the first all channels.
          
          Questionable:
          T3-1-V1 has channel 9-24 still high rate (threshold?) 
    	      No change from before.
          
    9:45  The Printer ran out of paper. If office supplies (printer paper,
          Ring Binders (of which also none are left) are running low
          please inform Melanie during the week, so the counting house 
          can be stocked.It is important, particular with one person shifts
          where one cannot leave the CH area for pickups.
    
    9:55 Voted to end-store; Star+Phenix had already done so.
    
    10:10 Mode switch back to gold - presumably preparing 
          for beam dump+refilling.
    
    10:15 MCR called - They will performe an aperture scan at STAR 
          within 10 minutes.
    
    10:24 Beam dumped. (Quench interlock).RUN 4741 Stopped.
    10:40. Turn magnets to standby in preparation for polarity Change.
           Looked at integral events for present running
    
           8 deg  A-pol (p-bar,K-)  
                  runs 4693 - 4741 :  Trigger 1 : 2,700 K
    			          Trigger 3 :   580 K
    			          Trigger 4 : 3,390 K
    	      It is noted that T3 was not operational for all of these runs.
           90 Deg 
    	      350A :  Trigger 1 : 765K
    	      350A :  Trigger 3 :   0K
    
    	      350B :  Trigger 1 : 1070K
    	      350B :  Trigger 3 :  260K
    
    11:13 Update the supermonitor to have autoprint for the Tof like detectors
          on the slat picture -not the bottom one.
    
    12:00 Magnet Polarity Changed. Went smoothly.
          The waweformgenerator for D3 has oscialting feedback, Current as well as
          Hall probe readout (check with Keithly web page) is stable.
          D1 676/B -2547 G  
          D2 552/B -3280 G
          D3 585/B -2945 G
          D4 527/B -3997 G
          D5 350/A Hall Keithley's ( 2087 Gauss)
          
    
    14:00 Despite a good ramp RHIC are having problems with cogging the beam. Though rates
          look reasonable it is not yet announched.See plot demonstaring the on-going acc work.
    
    14:15 Start run 4745. 
          Beam rates etc stable. The BB - Vertex distribution centered.
          Brahms rates ~2/3 of other exp as should be.
          Beam were manually cogged.
    
    
    15:30 Checked the Beam vertex. Was ~44 at start of store now ~ 46 cm Using the
          BB vertex dist in SuperMonitor
          ZDC rates ~ 65  singles Y: 6.5K B: 3.2K
    		      Ions B(19) Y(14)
          Yellow beam is decaying quite rapidly 20->14 in ~70 minutes.
    
    15:50 During this run MTPC2 readout has given errors multiple times  (6 so far) , where
          it has not reported the correct number of events. The sync event mechanism should
          take care of this, but please watch.
    
    --------------------------------------------------------------------------------
    
    Date
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b30 : Sat Sep 15 2001 - 23:05:41 EDT